Given the demand function Pd=25- 2Q and the supply function Ps= 2Q+1.Asssuming
pure competition, find the consumers’ surplus, producers’ surplus and interpret the total
of consumers’ surplus and producers’ surplus
1
Expert's answer
2013-08-06T08:23:01-0400
Consumers surplus calculated as Pd-M=25-2Q-M. Here M is a market price for a good.
Producers surplus equals M-Ps=M-2Q-1. It occurs when market price is greater than minimal price producer is willing to sell his good
for.
Total of producers and consumers surplus equals 25-2Q-M+M-2Q-1=24-4Qit's a difference between maximum price consumers are willing to pay and
minimum price producers are willing to sell for. As it reaches value of 0 (when
Q=6) market reaches its equilibrium.